Operation Summary for August 1, 2011 - August 3, 2011 |
Monday Aug 1 The day shift began with the Tevatron (TEV) off for cryo systems and vacuum systems recovery, with the Antiproton Source (Pbar) off, with the Recycler off, with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBFT-1008 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments off and conducting maintenance. During the day shift, FESS personnel reported at 9:43 AM that they had completed cleaning the TEV sector A3 pond water strainer. A Booster RF station (BRF16) tripped off at 9:15 AM for the first of three times in four minutes; it reset. Pbar personnel reported at 10:35 AM that they had successfully sent beam to Pbar and tested stacking. The Linac low energy RF station gallery air conditioning (AC) unit stopped working at 12:03 AM; operators set up fans. Operations turned off the Linac beam at 1:18 PM for AC work; FESS personnel completed their AC work and Linac resumed sending beam at 2:24 PM. At 2:58 PM, BRF17 & BRF18 tripped off on a ground fault indication; an RF person investigated, disconnected BRF18 from its power supply, and had it repaired and back online by 3:31 PM. During the evening shift, FESS personnel reported that the TEV sector A2 sump pump had been repaired. The Booster RF West Anode Power Supply (WAPS) tripped off at 5:02 PM for the first of 11 times in about 6.75 hours; it reset. Operations established beam to T-1008 at 5:47 PM. Operators tuned the MI accelerator. Tuesday Aug 2 The midnight shift began with the TEV off for cryo and vacuum systems recovery, and with the NuMI, MiniBooNE, and T-1008 experiments taking beam. During the midnight shift, the Booster RF WAPS tripped off at 12:48 AM for the first of 13 times in about 7.5 hours; it reset. At 6:15 AM, Cryo System personnel from CHL reported that they were having problems with their cold box; they began repairs. Operations disabled beam to T-1008 at 6:36 AM. Operators prepared the TEV Transfer Hall for access at 6:46 AM. The Neutron Therapy Facility (NTF) began treating patients at 7:27 AM. Operations disabled Booster beam at 7:54 AM so that RF personnel could investigate and repair a glycol leak in the WAPS. During the day shift, operators prepared the Switchyard enclosures for access at 8:21 AM to work on vacuum problems. Booster RF personnel reported at 8:43 AM that they had completed their leak repair on the WAPS. Operators prepared MI-31 and the Pelletron for access at 8:57 AM. At 9:21 AM, BRF16 tripped off for the first of two times in six minutes; it reset. Operations began sending beam to Pbar and Pbar began stacking at 11 AM. At 12:16 PM, BRF16 tripped off for the first of three times in 50 minutes; it reset. FESS personnel reported that they had cleaned the TEV sector F3 pond water strainer at 1:08 PM. NTF completed treating its last patient at 1:34 PM. During the evening shift, BRF15 & 16 tripped off; they reset. The Booster RF WAPS tripped off at 5:03 PM for the first of 10 times in about 5.5 hours; it reset. Operators turned on the TEV Separators at 8:21 PM. Operations established beam to T-1008 at 8:41 PM. At 10:05 PM, a lightning strike took down the Meson M01 and M02 Electrical Safety System (ESS) and dropped the enclosure interlocks; operators called in interlock technicians to investigate. Operators prepared M01 and M02 for access at 11:37 PM. Wednesday Aug 3 The midnight shift began with the TEV off for cryo and vacuum systems recovery, with Pbar stacking (33.7E10), with the NuMI and MiniBooNE experiments taking beam, and with T-1008 off due to electrical safety system work in M01 and M02. During the midnight shift, the Booster RF WAPS tripped off at 12:34 AM for the first of 3 times in about 2 hours; it reset. T-1008 decided to have their beam disabled at 12:51 AM due to the safety system problems; this ended their current run at Fermilab. Operators reported at 12:55 AM that a number of Pbar Debuncher stochastic cooling amplifiers tripped off due to rising cryo system temperatures. The Cryo Coordinator reported that the Pbar cryo system was dependent on nitrogen from the TEV; when the helium wave, presently running through the TEV, reaches TEV sector F3, sometime later today, things will get better. Interlock techs reported at 1:34 AM that they had repaired the ESS in M01 and M02. At 3:51 AM, a Linac quadrupole magnet (L:QTM2C) power supply tripped off; operators investigated and found the power supply had failed. Operators called in a Linac tech. He had the power supply back online by 4:51 AM. At 6:33 AM, a Booster ion pump (B:IPMP02) tripped off and took down B:MP02 for the first of three times in 20 minutes. Booster personnel investigated and made some changes. Operators turned off TEV separators at 7:04 AM and prepared the TEV A-E enclosures for access. The Booster RF WAPS tripped off at 7:46 AM for the first of 2 times in about 2 minutes; it reset. A Pbar stochastic cooling person began work on the Recycler stochastic cooling problems. The Plans for Wednesday The plans for today are to continue with TEV cryo and vacuum systems recovery. The Cryo Coordinator expects the TEV cryo system to be ready by Friday morning (8/5/11). The Run Coordinator will allow Booster and MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 08/03/2011 email Fermilab |
| Security, Privacy, Legal | |